rvc (rvc):

Please help me :) If x, y, z are in A.P. and tan−1x, tan−1y and tan−1z are also in A.P., then 1. 2x = 3y = 6z 2. 6x = 3y = 2z 3. 6x = 4y = 3z 4. x = y = z

Miracrown (miracrown):

Ahoy Matey! What are the directions for the problem?

Miracrown (miracrown):

Just to make sure, what is A.P. ?

rvc (rvc):

arithematic progression

Miracrown (miracrown):

Ok, thx matey! My brain was goin' nuts there ok so do you remember the definition of an Arithmetic Progression?

rvc (rvc):

well how do i explain now :)

Miracrown (miracrown):

Let me write an example so it can be clearer as as sometimes definitions are hard to understand... 2,5,8 <--- these numbers are in arithmetic progression, as the difference between them (consecutives) are always equal - let's check it 5-2 = 3 8-5 = 3 And that is called the common difference If you have 3 numbers in A.P. you can write a relation between them as they have the common difference, in our case we do have x, y and z, what can you say about them if they are in A.P. ?

rvc (rvc):

well tbh i know that thats for explaining

rvc (rvc):

\[t _{n}=a+(n-1)d\]

rvc (rvc):

a=first term d= common difference n= term

Miracrown (miracrown):

Z – y Y – x we do not know what is the common difference equal to Z – y = y – x <---- but what we do know is that it is equal to both expressions And we can work a little bit with it. Let’s isolate y, then add x at both sides so we can isolate 2y And we get x + z = 2y

rvc (rvc):

\[\tan^{-1} x ,\tan^{-1} y,\tan^{-1} z\]

rvc (rvc):

wait i can see question marks what did u type

Miracrown (miracrown):

Refresh page matey its a odd glitch

Miracrown (miracrown):

\(\color{blue}{\text{Originally Posted by}}\) @rvc \[t _{n}=a+(n-1)d\] \(\color{blue}{\text{End of Quote}}\) That is the formula for the general term but we do not need to use it as we just have 3 terms and we found the relation for them using the basic principles

rvc (rvc):

yes i know i just made u sure that i know A.P :)

Miracrown (miracrown):

Hehe matey, I know, I know! And now would you please write the same relation for tan^-1x , tan^-1y and tan^-1z

rvc (rvc):

well those are inverse functions okay:)

rvc (rvc):

ah im lagging sorry

Miracrown (miracrown):

Same here! The lag is real. ;-; Anyhow, we do have tan^-1 x tan^-1 y and tan^-1 z, we can write the same expression for those terms... would ye please write it?

Miracrown (miracrown):

We just need to start with the double of the middle term, what is the middle term?

rvc (rvc):

\[\tan^{-1} y-\tan^{-1} x ,\tan^{-1} z-\tan^{-1} y\] did u mean this :)

Miracrown (miracrown):

Ye! (^: nd then solve it for 2tan^-1(y)

rvc (rvc):

\[2\tan^{-1} y=\tan^{-1} x+\tan^{-1} z\]

Miracrown (miracrown):

That's absolutely r8 matey!!! Ye ho

rvc (rvc):

may i know what's matey?

Miracrown (miracrown):

Ehehehe, it's a word often used by pirates to address friends (;

rvc (rvc):

cool :)

Miracrown (miracrown):

Ye ho! Also remember please ALWAYS!: if 3 terms are in A.P. then the double of the middle term is equal to the sum of the other 2 terms and you can see it in both expression 2 times the middle term equal to the sum of the other terms and the only solution for this system of equation is when all thre are equal to each other in other words> x = y = z

Miracrown (miracrown):

..And... let's verify it, if x = y and z = y would you please plug it?

rvc (rvc):

hmm.. so what should i do?

Miracrown (miracrown):

You just need to verify it as x = y and z = y, plug instead of x y and instead of z y to verify we got it correctly

rvc (rvc):

oh

rvc (rvc):

oh the RHS =LHS

Miracrown (miracrown):

you just need to verify the solution as x = y you plug instead of x just y and the same for z so you can see that it is correct right hand = left hand, yes matey

rvc (rvc):

so the correct ans is x=y=z @Miracrown thank you so much :)

rvc (rvc):

thank you for giving your time @Miracrown You helped me a lot :)

Miracrown (miracrown):

Ye, that's right on the bat! And no worries matey! (^: